OpenStudy (anonymous):

A falling stone is at a certain instant 100 feet above the ground. Two seconds later it is only 20 feet above the ground. If it was thrown down with an initial speed of 7 feet per second, from what height was it thrown?

OpenStudy (anonymous):

This is a calculus problem not phsyics

OpenStudy (anonymous):

. I am assuming that you drop the stone from rest, not throw it down or up. ok falling objects accelerate to the ground at the constant rate 32 ft/s^2 drop object from some height at time 0. After t secs, since the object is dropped from rest, the speed of the object is v= 32t ft/s distance object has travelled at time t from the spot you dropped it is s=16t^2 ft (how did i get this? distance is the area under the speed graph. so i integrated 32t) a. say the object is dropped from height h. then when it is 100 ft above ground, it has travelled h-100 ft from dropping point. how much time passed? h-100 = 16t^2 ----> t= 0.25*√(h-100) 2 seconds later it is only 16 ft above ground (so it has travelled h-16 ft) h-16 = 16(t+2)^2 h-16 = 16*(0.25*√(h-100)+2)^2 this is an equation in h..... solve it! if you're allowed to use a graphing calculator then its easy.

OpenStudy (anonymous):

The answers are 6479/64

OpenStudy (anonymous):

6415/64

OpenStudy (anonymous):

6607/64

OpenStudy (anonymous):

let H(t) = yo -7(t)-16t^{2} H(c) = 100 = yo -7(c) -16c^{2} H(c+2) = 20 = yo -7(c+2) - 16(c+2)^{2} eq1) 20 = yo -7c -14 - 16(c^{2}+4c + 4) yo = 20 +14 +7c + 16c^{2} +64c +64 eq2) yo = 100 + 7c +16c^{2} solve eq1=eq2 100 + 7c + 16^{2} = 34 + 7c + 16c^{2} +64c+64 2 = 64c c = 1/32 now H(1/32) = 100=yo - 7(1/32) - 16(1/34)^{2}

OpenStudy (anonymous):

yo = 100 15/64 = 6415/64

OpenStudy (anonymous):

that part up there is -16(1/32)^{2

OpenStudy (anonymous):

Thankh